determine number of oxydation of the following CAPITALS elements

a) O3
b) h3PO4
c)MNo3
d)C2o4 -2
1
Expert's answer
2018-08-03T05:07:15-0400
a) The number of oxydation for O is zero, because O3 is a simple substance.
b) The number of oxydation for H is +1, O is -2. The number of oxydation for P we can find by solving equation:
(+1)*3+x+(-2)*4=0
x=+5
c) For Mn we have:
x+(-2)*3=0
x=+6
d) For C we have:
2x+(-2)*4=-2
x=+3